ques: the speed of sound in oxygen at a certain temp is 460 ms -1 .the speed of sound in he at same temp will be (assume both gases to be ideal) a)460 ms -1 c)650 ms-1 b)500 ms -1 d)330 ms-1 ques: an athele in the olympics games covers a distance of 100 m in 10 s.his kinetic energy can be estimated to be in the range a)200j-500j c)2x10 5j -3x10 5j b)20,000j-50,000j d)2,000j-5,000j thanks

ques:the speed of sound in oxygen at a certain temp is 460 ms-1 .the speed of sound in he at same temp will be (assume both gases to be ideal)


a)460 ms -1                        c)650 ms-1


b)500 ms  -1                       d)330 ms-1


ques:an athele in the olympics games covers a distance of 100 m in 10 s.his kinetic energy can be estimated to be in the range


a)200j-500j                       c)2x105j-3x105j


b)20,000j-50,000j               d)2,000j-5,000j


thanks

Grade:12

1 Answers

Askiitians Expert Soumyajit IIT-Kharagpur
14 years ago

Dear Charu Gupta,

Ans;

Qsn no 1

the speed of sound in a gas is V=(γRT/M)^(.5)

So if T is constant, We have V∞(γ/M)^.5

for O2 γ=7/5, M=32

For He γ=5/3, M=4;

hence the velocity in He is=velocity in O2×√[(5/3 ÷4)(32÷7/5)]

=1419.6 m/s

hence none of the options is correct.

Question 2

The kinetic energy is mass dependent which you have not given and hence the KE can't be calculated using your data.

 

Please feel free to post as many doubts on our discussion forum as you can. If you find any question
Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We
are all IITians and here to help you in your IIT JEE preparation.


All the best Charu Gupta !!!

 


Regards,

Askiitians Experts
Soumyajit Das IIT Kharagpur